# Ultrafast generation of coherent soft-shear phonons in halide perovskites via anisotropic photostriction

### v1

@@ I Introduction
Terahertz (THz) and sub-THz coherent acoustic phonons show great potential for manipulations involving photons [1, 2, 3, 4, 5, 6], electrical currents [7, 8] and fields [9], magnons [10, 11], as well as plasmons [12]. Experiments have demonstrated that high frequency coherent phonons can be used for the generation of THz electromagnetic waves [13, 14]. Due to their nanometer wavelength, sub-THz phonons are applied for nano-imaging [15]. Ultrafast spectroscopy of high-frequency coherent acoustic phonons employs the excitation and detection using short femtosecond laser pulses in pump-probe techniques, representing a powerful method to investigate the lattice dynamics in a wide spectrum of materials [16, 17]. Typically, the generation of coherent phonons exploits the excitation of ultrafast stress associated with the thermoelastic coupling in metals [16, 17, 18, 19] or with the electronic deformation potential in semiconductors [16, 20, 21]. In most settings, the phonons are generated with polarization vector $\mathbf{U}$ parallel to the propagation direction and correspondingly to the phonon wave vector $\mathbf{Q}$, i.e. longitudinal acoustic (LA) phonons are provided. However, experiments with coherent phonons, whose polarization is perpendicular to $\mathbf{Q}$, i.e. transverse acoustic (TA, shear) phonons, have remained in high demand. Indeed, TA phonons possess a smaller sound velocity compared to LA phonons and, correspondingly, a smaller wavelength for the same frequency, of great advantage in imaging. Further, TA phonons posses two independent polarization components and could be used to manipulate the spin properties of charge carriers by chiral acoustic waves [22].
@@
Generation of TA phonons requires a shear perturbation of the crystalline lattice. The most common method to induce shear strain in a material is to excite a crystalline surface of low symmetry by an ultrafast optical pulse  [23, 24, 25]. In this case, even isotropic compressive stress results in both compressive and shear dynamical strain and leads to the generation of quasi-LA and quasi-TA phonons with wave vectors $\mathbf{Q}$ perpendicular to the optically excited surface. Also, the piezoelectric mechanism of phonon generation may have significant importance [26, 27]. Usually, the amplitude of the generated TA phonons is much smaller than that of the LA phonons. Only recently, a few studies have reported amplitudes of the generated TA phonons comparable with those of the LA phonons  [28, 29, 30], but the mechanisms of efficient generation of shear acoustic waves are still under debate. Here, material systems such as the perovskites exhibiting a variety of structural phase transitions, provide a rich testbed for exploring new scenarios of TA phonon generation. Furthermore, semiconductor perovskites are particularly interesting for exploiting photoinduced mechanisms as the interaction of optically excited carriers with the crystal are dominant here.
@@
In recent years, the interest in perovskite semiconductors has grown rapidly due to their success in photovoltaic applications [31]. Inorganic lead-free double perovskites are particularly relevant here as non-toxic and stable material platform, but they so far suffer from inefficient charge generation and transport, for which the responsible mechanisms are yet not well understood [32, 33]. Among them are complex phenomena associated with structural phase transitions (cubic to tetragonal) [34] and strong electron-phonon interactions (polaron effects) [32, 35], which can significantly influence the mobility of charge carriers [36]. In this respect, double perovskite semiconductors are particularly exciting for ultrafast acoustics. First, the elastic constants are about three times smaller as compared to conventional semiconductors such as GaAs [37]. A soft crystal lattice allows one to achieve a significantly larger deformation for the same magnitude of stress. Second, despite the absence of the piezoelectric contribution leading to a strong renormalization of the photoinduced stress in ferroelectric perovskites [38], in lead halide perovskites an analogous renormalization due to structural phase transition is present [39]. Moreover, the formation of exciton polarons, which distort the lattice after pulsed optical excitation, is not captured by the “deformation potential” mechanism of coherent phonon generation. Finally, the ease of their synthesis as well as the rapid development of growth technology of perovskite semiconductors underline the potential for deposition of thin films on top of other materials which could subsequently lead to the realization of optically driven phononic transducers [40].
@@ II Ultrafast optical response in $\mathbf{Cs_{2}AgBiBr_{6}}$
Single crystals of Cs2AgBiBr6 were grown from a mixture of CsBr, AgBr and BiBr in a 2:1:1 molar ratio in 48% wt HBr [41, 42]. The reaction flask was heated to 110C for complete dissolution of the metal salts, and cooled down at a rate of 1C/h to room temperature (see Methods). The simple cubic crystal structure in ambient conditions was proven by powder X-ray diffraction (see Supplementary Information section S1). All of the crystal facets are associated with (111) or equivalent crystallographic planes as shown in Fig. 1(a). Note that in the tetragonal phase these planes have the Miller indices (011) and equivalent.
@@
The optical properties of the investigated sample are summarized in Fig. 1(b). The low temperature reflectivity spectrum (blue line) shows a strong feature centered at about 2.822 eV, which is associated with the exciton resonance corresponding to the direct interband optical transition at the $\Gamma$-point [35]. The green and red lines in Fig. 1(b) show the transmission and photoluminescence (PL) spectra with a sharp absorption edge at about 2.4 eV and a broad PL peak at about 2.0 eV, respectively, in agreement with previous studies [35, 36]. Note that these values cannot be directly attributed to the indirect band gap due to the strong electron-phonon coupling with a Huang-Rhys factor in the order of 10 [35].
For investigation of the optically generated strain pulses, we use the well established method of time-domain Brillouin spectroscopy [27]. As shown in Fig. 1(c), the pump pulse with a photon energy above the absorption edge generates photoexcited carriers in the thin surface area of the crystal, which leads to the formation of a stress driving force $\mathbf{\Phi}$ and to subsequent propagation of a strain pulse away from the surface along the $\mathbf{z}^{\prime}$ direction, i.e. along the gradient of the photoexcited carrier density $\nabla\rho_{X}$. In isotropic or cubic crystals with high symmetry crystallographic directions, $\mathbf{\Phi}\|\mathbf{z}^{\prime}$ so that only longitudinal strain pulses are generated. However, as will be shown below, anisotropic expansion and, in particular, simultaneous expansion and compression along different directions can also induce a force with a direction different from $\mathbf{z}^{\prime}$, which subsequently leads to the generation of shear strain pulses (TA pulse in Fig. 1(c)).
@@
Surprisingly, at the low temperature of $T=5$ K, when the crystal is in the tetragonal phase, the signal comprises two frequencies $f_{i}$ of about 10 and 23 GHz, see Fig. 2(c). At temperatures above $T_{c}\approx 122$ K, the crystal structure transforms into the cubic phase [34]. Here, only the component with the higher frequency persists, as shown in the insets of Fig. 2(a) and Fig. 2(c) for the temperature of 150 K. The normalized fast Fourier transform (FFT) spectra of the oscillatory signals for HF excitation shown in Fig. 2(a) are presented in Fig. 2(c). The frequencies are independent of the photon energy of the pump pulses, but increase linearly with increase of the probe photon energy $h\nu_{\rm probe}$, see Fig. 2(d).
This behavior is observed in time-domain Brillouin scattering on acoustic phonons, i.e. when the probe pulse is scattered from the pump-induced strain pulse propagating away from the surface with the speed of sound $V_{i}$ [27]. In this case, the oscillation frequency is determined as $f_{i}=2n_{r}\nu_{\rm probe}V_{i}/c$, where $c$ is the speed of light and $n_{r}$ is the refractive index in Cs2AgBiBr6. Taking the refractive index at room temperature $n_{r}=2$ from Ref. [43], we fit the data in Fig. 2(d) with a linear function and estimate the speed of sound for the LA and one of the TA modes as $V_{\rm LA}=2900$ m/s and $V_{\rm TA_{y}}=1400$ m/s at $T=5$ K, which are close to the data evaluated by $cw$ Brillouin light scattering [44].
@@ III Optical generation of shear strain in tetragonal phase
The stress tensor contains two main contributions. The first one is given by the linear expansion of the lattice induced by heat [16] which is proportional to the increase of the lattice temperature $\delta T$:
@@
Here, $\alpha_{i}$ is the photostriction coefficient, i.e. the change of lattice constant by the photogenerated electron-hole pairs. In the literature, the stress due to the presence of carriers is called either “deformation potential” or “photostriction”. We use the second term to highlight that the deformation potential is not the only mechanism contributing to the phenomenon described by Eq. (5). As shown recently for lead halide perovskites in Ref. [39], the deformation potential alone is insufficient to describe the effect of the charge concentration on stress, not even qualitatively.
In GaAs the strain caused by lattice heating as described by Eq. (4) is estimated to be $\sim 3\%$ from the contribution given by Eq. (5[45]. In MAPbI3, the photostriction coefficient is estimated [46] to be at least two orders of magnitude larger than in GaAs. Therefore, it appears to be consequent to expect that this contribution is the dominating one in (double) perovskites.
@@
where $\zeta$ is the absorption length with $\bf r$ being the coordinate. In case of isotropic thermal expansion and photostriction the driving force is directed along the surface normal $\bm{\Phi}\|{\bf n}$ and only LA phonons can be generated. However, for materials of low symmetry, the linear expansion coefficient may be different for different crystallographic directions [47]. The thermal coefficients for the linear expansion in Cs2AgBiBr6 are almost equal in magnitude, but have different signs for the directions perpendicular to and along the $[001]$ crystal axis [34]. This behavior is also typical for tetragonal phase perovskites: in $\beta$-MAPbI3 (space group #140, $I4/mcm$), $\beta_{\perp c}\approx-\beta_{||c}$ [48].
To our knowledge, neither measurements nor calculations of the photostriction coefficients for Cs2AgBiBr6 are available. However, calculations for $\beta$-CsPbI3 show that $\alpha$ has different signs perpendicular to and along the $c$-axis ($\alpha_{\perp c}/\alpha_{||c}\approx-0.6$[39], similar to the temperature expansion coefficient. In CsPbI3, this large anisotropy has been qualitatively explained by accounting for the cubic-to-tetragonal phase transition [38, 39]. Recently, it also has been demonstrated that one of the main contributions to the electron-phonon interaction in CsPbI3 is the interaction with the LO phonon modes associated with phase transitions [49]. Therefore, we expect that a similarly anisotropic photostriction is present in double perovskites showing cubic to tetragonal phase transition and strong polaron effects.
@@
The detected signal arises from the reflection of the probe pulse at the strain wave. Its strength is given by the strain-induced variation of the dielectric constant $\Delta\hat{\varepsilon}$ [24, 50] which in turn is proportional to the strain tensor $\hat{\epsilon}$ [51, 52, 53]. As shown in Sec. S4 of the Supplementary Information, the strain tensor for different modes is proportional not only to the amplitude $\sim{\bm{\Phi}}\cdot{\bf U}^{s}$, but also to $\epsilon^{s}_{ij}=\frac{1}{2}(U_{i}^{s}Q_{j}+U_{i}^{s}Q_{i})$.
@@ IV Temperature dependence: tetragonal vs cubic phase
The difference in the temperature at which the phase transition occurs in the $cw$ and time-domain BLS is attributed to local heating of the sample by the laser, which is most pronounced for HF excitation for which the integrated power is very high (2 kW/cm2) so that the actual crystal temperature within the excitation spot is about 40 K higher than the bath temperature. For $cw$ BLS the excitation power is even higher (100 kW/cm2) but the photon energy of 2.287 eV is below the direct band gap and the corresponding increase of temperature is only about 10 K. This is in contrast to the negligible heating observed in the LF pump-probe data with a time-integrated power of about $20$ W/cm2, for which the phase transition temperature corresponds to the value of 122 K reported in literature [34].
@@ Discussion
The weak temperature dependence of the photogenerated strain evidences the non-thermal nature of the effect, i.e., the direct generation of strain via carrier-phonon interaction accomplished due to the anisotropic photostriction. Indeed, the heat capacity has a strong nonlinear $T$-dependence, which possesses a sharp increase below 70 K and a slow variation at higher temperatures, while the thermal expansion coefficient does not depend on temperature [34].
Using the temperature dependence of the heat capacity from Ref. [34] and the absorption coefficient $\alpha=1.5\times 10^{5}$ cm-1 from Ref. [43], we can estimate the increase of the temperature $\delta T$ after excitation with the maximum fluence of $\Psi_{\rm pump}=0.7$ mJ/cm2. For instance, at $T=5$ K the optical excitation leads to $\delta T\approx 47$ K, while for $T=100$ K the temperature increase is significantly reduced to $\delta T(100\text{\leavevmode\nobreak\ K})=23$ K. Using the linear temperature expansion coefficient of $|\beta_{i}|\approx 10^{-5}$ K-1, we estimate the amplitude of the strain field to be in the order of $10^{-3}$. Note that $\delta T(5\text{\leavevmode\nobreak\ K})/\delta T(100\text{\leavevmode\nobreak\ K})\sim 2$ means that the temperature effect on the strain generation should result in a strong change (at least by a factor of two) of the signal amplitude when the temperature of the sample changes from 5 to 100 K. The weak dependence of the photoinduced strain on temperature in this range confirms that the main effect is not associated to the temperature-based mechanism (see Eq. (4)) and occurs mainly due to the photostriction, given by Eq. (5). Assuming that the photostriction mechanism is at least as strong as that through the temperature change, we can estimate the photostriction coefficient $\alpha_{i}$. Using the exciton density $\rho_{X}(0)\approx 5\times 10^{19}$cm-3 near the surface, we obtain the photostriction coefficient $|\alpha_{i}|\sim 10^{-23}$ cm3.
We stress that the anisotropic photostriction is directly related to the strong interaction of electrons with optical phonons in combination with the crystallographic phase transition. Therefore, the excitation of soft TA coherent phonons could be due to interaction with the specific optical phonon eigenmodes responsible for the tetragonal to cubic phase transition. Particularly such optical phonons posses a strong interaction with charge carriers [49]. Further studies of the microscopic origin of the anisotropic photostriction could allow one to adjust the parameters of the material for efficient photoinduced generation of shear strain and realization of perovskite-based phononic transducers, including those for chiral phonons.

### v2

# Efficient launch of shear phonons in photostrictive halide perovskites

@@ I Introduction
Terahertz (THz) and sub-THz coherent acoustic phonons show great potential for manipulations involving photons [Fainstein-2024, Brueggemann2012, Ortiz-21, Chern-2004, Matsuda-2002, Baldini-2019], electrical currents [Sun-2004, Kent-2012] and fields [Ruello-2023], magnons [Sherbakov-2010, Reppert-2020], as well as plasmons [Temnov-2012]. Experiments have demonstrated that high frequency coherent phonons can be used for the generation of THz electromagnetic waves [Armstrong-2009, Rongione-2023]. Due to their nanometer wavelength, sub-THz phonons are applied for nano-imaging [Audoin-2023]. Ultrafast spectroscopy of high-frequency coherent acoustic phonons employs the excitation and detection using short femtosecond laser pulses in pump-probe techniques, representing a powerful method to investigate the lattice dynamics in a wide spectrum of materials [Thomsen1986, Matsuda2015]. Typically, the generation of coherent phonons exploits the excitation of ultrafast stress associated with the thermoelastic coupling in metals [Thomsen1986, Matsuda2015, Maris-1994, Wright-1994] or with the electronic deformation potential in semiconductors [Thomsen1986, Young-2012, Lamitre-2012]. In most settings, the phonons are generated with polarization vector $\mathbf{U}$ parallel to the propagation direction and correspondingly to the phonon wave vector $\mathbf{Q}$, i.e. longitudinal acoustic (LA) phonons are provided. However, experiments with coherent phonons, whose polarization is perpendicular to $\mathbf{Q}$, i.e. transverse acoustic (TA, shear) phonons, have remained in high demand. Indeed, TA phonons possess a smaller sound velocity compared to LA phonons and, correspondingly, a smaller wavelength for the same frequency, of great advantage in imaging. Further, TA phonons posses two independent polarization components and could be used to manipulate the spin properties of charge carriers by chiral acoustic waves [Korenev-2016].
@@
Generation of TA phonons requires a shear perturbation of the crystalline lattice. The most common method to induce shear strain in a material is to excite a crystalline surface of low symmetry by an ultrafast optical pulse  [Gusev-2007, Matsuda2008, Kent-2017]. In this case, even isotropic compressive stress results in both compressive and shear dynamical strain and leads to the generation of quasi-LA and quasi-TA phonons with wave vectors $\mathbf{Q}$ perpendicular to the optically excited surface. Also, the piezoelectric mechanism of phonon generation may have significant importance [Sun-2012, RUELLO201521]. Usually, the amplitude of the generated TA phonons is much smaller than that of the LA phonons. Only recently, a few studies have reported amplitudes of the generated TA phonons comparable with those of the LA phonons  [Lejman-2014, Guo-2017, Mante-2018], but the mechanisms of efficient generation of shear acoustic waves are still under debate. Here, material systems such as the perovskites exhibiting a variety of structural phase transitions, provide a rich testbed for exploring new scenarios of TA phonon generation. Furthermore, semiconductor perovskites are particularly interesting for exploiting photoinduced mechanisms as the interaction of optically excited carriers with the crystal are dominant here.
@@
In recent years, the interest in perovskite semiconductors has grown rapidly due to their success in photovoltaic applications [Roadmap-2021]. Inorganic lead-free double perovskites are particularly relevant here as non-toxic and stable material platform, but they so far suffer from inefficient charge generation and transport, for which the responsible mechanisms are yet not well understood [Wu-2021, Tress-2022]. Among them are complex phenomena associated with structural phase transitions (cubic to tetragonal) [Schade2019] and strong electron-phonon interactions (polaron effects) [Wu-2021, Baranowski-2019], which can significantly influence the mobility of charge carriers [Herz-THz-2021]. In this respect, double perovskite semiconductors are particularly exciting for ultrafast acoustics. First, the elastic constants are about three times smaller as compared to conventional semiconductors such as GaAs [Even-2016]. A soft crystal lattice allows one to achieve a significantly larger deformation for the same magnitude of stress. Second, despite the absence of the piezoelectric contribution leading to a strong renormalization of the photoinduced stress in ferroelectric perovskites [Paillard2017], in lead halide perovskites an analogous renormalization due to structural phase transition is present [Paillard2023]. Moreover, the formation of exciton polarons, which distort the lattice after pulsed optical excitation, is not captured by the “deformation potential” mechanism of coherent phonon generation. Finally, the ease of their synthesis as well as the rapid development of growth technology of perovskite semiconductors underline the potential for deposition of thin films on top of other materials which could subsequently lead to the realization of optically driven phononic transducers [Du2021].
@@ II Ultrafast optical response in $\mathbf{Cs_{2}AgBiBr_{6}}$
Single crystals of Cs2AgBiBr6 were grown from a mixture of CsBr, AgBr and BiBr in a 2:1:1 molar ratio in 48% wt HBr [Adam-2016, Armer-2021]. The reaction flask was heated to 110C for complete dissolution of the metal salts, and cooled down at a rate of 1C/h to room temperature (see Methods). The simple cubic crystal structure in ambient conditions was proven by powder X-ray diffraction (see Supplementary Information section S1). All of the crystal facets are associated with (111) or equivalent crystallographic planes as shown in Fig. 1(a). Note that in the tetragonal phase these planes have the Miller indices (011) and equivalent.
@@
The optical properties of the investigated sample are summarized in Fig. 1(b). The low temperature reflectivity spectrum (blue line) shows a strong feature centered at about 2.822 eV, which is associated with the exciton resonance corresponding to the direct interband optical transition at the $\Gamma$-point [Baranowski-2019]. The green and red lines in Fig. 1(b) show the transmission and photoluminescence (PL) spectra with a sharp absorption edge at about 2.4 eV and a broad PL peak at about 2.0 eV, respectively, in agreement with previous studies [Baranowski-2019, Herz-THz-2021]. Note that these values cannot be directly attributed to the indirect band gap due to the strong electron-phonon coupling with a Huang-Rhys factor in the order of 10 [Baranowski-2019].
For investigation of the optically generated strain pulses, we use the well established method of time-domain Brillouin spectroscopy [RUELLO201521]. As shown in Fig. 1(c), the pump pulse with a photon energy above the absorption edge generates photoexcited carriers in the thin surface area of the crystal, which leads to the formation of a stress driving force $\mathbf{\Phi}$ and to subsequent propagation of a strain pulse away from the surface along the $\mathbf{z}^{\prime}$ direction, i.e. along the gradient of the photoexcited carrier density $\nabla\rho_{X}$. In isotropic or cubic crystals with high symmetry crystallographic directions, $\mathbf{\Phi}\|\mathbf{z}^{\prime}$ so that only longitudinal strain pulses are generated. However, as will be shown below, anisotropic expansion and, in particular, simultaneous expansion and compression along different directions can also induce a force with a direction different from $\mathbf{z}^{\prime}$, which subsequently leads to the generation of shear strain pulses (TA pulse in Fig. 1(c)).
@@
Surprisingly, at the low temperature of $T=5$ K, when the crystal is in the tetragonal phase, the signal comprises two frequencies $f_{i}$ of about 10 and 23 GHz, see Fig. 2(c). At temperatures above $T_{c}\approx 122$ K, the crystal structure transforms into the cubic phase [Schade2019]. Here, only the component with the higher frequency persists, as shown in the insets of Fig. 2(a) and Fig. 2(c) for the temperature of 150 K. The normalized fast Fourier transform (FFT) spectra of the oscillatory signals for HF excitation shown in Fig. 2(a) are presented in Fig. 2(c). The frequencies are independent of the photon energy of the pump pulses, but increase linearly with increase of the probe photon energy $h\nu_{\rm probe}$, see Fig. 2(d).
This behavior is observed in time-domain Brillouin scattering on acoustic phonons, i.e. when the probe pulse is scattered from the pump-induced strain pulse propagating away from the surface with the speed of sound $V_{i}$ [RUELLO201521]. In this case, the oscillation frequency is determined as $f_{i}=2n_{r}\nu_{\rm probe}V_{i}/c$, where $c$ is the speed of light and $n_{r}$ is the refractive index in Cs2AgBiBr6. Taking the refractive index at room temperature $n_{r}=2$ from Ref. [Huygen2021], we fit the data in Fig. 2(d) with a linear function and estimate the speed of sound for the LA and one of the TA modes as $V_{\rm LA}=2900$ m/s and $V_{\rm TA_{y}}=1400$ m/s at $T=5$ K, which are close to the data evaluated by $cw$ Brillouin light scattering [Lun2022].
@@ III Optical generation of shear strain in tetragonal phase
The stress tensor contains two main contributions. The first one is given by the linear expansion of the lattice induced by heat [Thomsen1986] which is proportional to the increase of the lattice temperature $\delta T$:
@@
Here, $\alpha_{i}$ is the photostriction coefficient, i.e. the change of lattice constant by the photogenerated electron-hole pairs. In the literature, the stress due to the presence of carriers is called either “deformation potential” or “photostriction”. We use the second term to highlight that the deformation potential is not the only mechanism contributing to the phenomenon described by Eq. (5). As shown recently for lead halide perovskites in Ref. [Paillard2023], the deformation potential alone is insufficient to describe the effect of the charge concentration on stress, not even qualitatively.
In GaAs the strain caused by lattice heating as described by Eq. (4) is estimated to be $\sim 3\%$ from the contribution given by Eq. (5[Matsuda2005]. In MAPbI3, the photostriction coefficient is estimated [Zhou2016] to be at least two orders of magnitude larger than in GaAs. Therefore, it appears to be consequent to expect that this contribution is the dominating one in (double) perovskites.
@@
where $\zeta$ is the absorption length with $\bf r$ being the coordinate. In case of isotropic thermal expansion and photostriction the driving force is directed along the surface normal $\bm{\Phi}\|{\bf n}$ and only LA phonons can be generated. However, for materials of low symmetry, the linear expansion coefficient may be different for different crystallographic directions [ITCBD]. The thermal coefficients for the linear expansion in Cs2AgBiBr6 are almost equal in magnitude, but have different signs for the directions perpendicular to and along the $[001]$ crystal axis [Schade2019]. This behavior is also typical for tetragonal phase perovskites: in $\beta$-MAPbI3 (space group #140, $I4/mcm$), $\beta_{\perp c}\approx-\beta_{||c}$ [Jacobsson2015].
To our knowledge, neither measurements nor calculations of the photostriction coefficients for Cs2AgBiBr6 are available. However, calculations for $\beta$-CsPbI3 show that $\alpha$ has different signs perpendicular to and along the $c$-axis ($\alpha_{\perp c}/\alpha_{||c}\approx-0.6$[Paillard2023], similar to the temperature expansion coefficient. In CsPbI3, this large anisotropy has been qualitatively explained by accounting for the cubic-to-tetragonal phase transition [Paillard2017, Paillard2023]. Recently, it also has been demonstrated that one of the main contributions to the electron-phonon interaction in CsPbI3 is the interaction with the LO phonon modes associated with phase transitions [Hoffman2023]. Therefore, we expect that a similarly anisotropic photostriction is present in double perovskites showing cubic to tetragonal phase transition and strong polaron effects.
@@
The detected signal arises from the reflection of the probe pulse at the strain wave. Its strength is given by the strain-induced variation of the dielectric constant $\Delta\hat{\varepsilon}$ [Matsuda2008, Subbaswamy1978] which in turn is proportional to the strain tensor $\hat{\epsilon}$ [Nelson70, Anastassakis74, Vacher72]. As shown in Sec. S4 of the Supplementary Information, the strain tensor for different modes is proportional not only to the amplitude $\sim{\bm{\Phi}}\cdot{\bf U}^{s}$, but also to $\epsilon^{s}_{ij}=\frac{1}{2}(U_{i}^{s}Q_{j}+U_{i}^{s}Q_{i})$.
@@ IV Temperature dependence: tetragonal vs cubic phase
The difference in the temperature at which the phase transition occurs in the $cw$ and time-domain BLS is attributed to local heating of the sample by the laser, which is most pronounced for HF excitation for which the integrated power is very high (2 kW/cm2) so that the actual crystal temperature within the excitation spot is about 40 K higher than the bath temperature. For $cw$ BLS the excitation power is even higher (100 kW/cm2) but the photon energy of 2.287 eV is below the direct band gap and the corresponding increase of temperature is only about 10 K. This is in contrast to the negligible heating observed in the LF pump-probe data with a time-integrated power of about $20$ W/cm2, for which the phase transition temperature corresponds to the value of 122 K reported in literature [Schade2019].
@@ Discussion
The weak temperature dependence of the photogenerated strain evidences the non-thermal nature of the effect, i.e., the direct generation of strain via carrier-phonon interaction accomplished due to the anisotropic photostriction. Indeed, the heat capacity has a strong nonlinear $T$-dependence, which possesses a sharp increase below 70 K and a slow variation at higher temperatures, while the thermal expansion coefficient does not depend on temperature [Schade2019].
Using the temperature dependence of the heat capacity from Ref. [Schade2019] and the absorption coefficient $\alpha=1.5\times 10^{5}$ cm-1 from Ref. [Huygen2021], we can estimate the increase of the temperature $\delta T$ after excitation with the maximum fluence of $\Psi_{\rm pump}=0.7$ mJ/cm2. For instance, at $T=5$ K the optical excitation leads to $\delta T\approx 47$ K, while for $T=100$ K the temperature increase is significantly reduced to $\delta T(100\text{\leavevmode\nobreak\ K})=23$ K. Using the linear temperature expansion coefficient of $|\beta_{i}|\approx 10^{-5}$ K-1, we estimate the amplitude of the strain field to be in the order of $10^{-3}$. Note that $\delta T(5\text{\leavevmode\nobreak\ K})/\delta T(100\text{\leavevmode\nobreak\ K})\sim 2$ means that the temperature effect on the strain generation should result in a strong change (at least by a factor of two) of the signal amplitude when the temperature of the sample changes from 5 to 100 K. The weak dependence of the photoinduced strain on temperature in this range confirms that the main effect is not associated to the temperature-based mechanism (see Eq. (4)) and occurs mainly due to the photostriction, given by Eq. (5). Assuming that the photostriction mechanism is at least as strong as that through the temperature change, we can estimate the photostriction coefficient $\alpha_{i}$. Using the exciton density $\rho_{X}(0)\approx 5\times 10^{19}$cm-3 near the surface, we obtain the photostriction coefficient $|\alpha_{i}|\sim 10^{-23}$ cm3.
We stress that the anisotropic photostriction is directly related to the strong interaction of electrons with optical phonons in combination with the crystallographic phase transition. Therefore, the excitation of soft TA coherent phonons could be due to interaction with the specific optical phonon eigenmodes responsible for the tetragonal to cubic phase transition. Particularly such optical phonons posses a strong interaction with charge carriers [Hoffman2023]. Further studies of the microscopic origin of the anisotropic photostriction could allow one to adjust the parameters of the material for efficient photoinduced generation of shear strain and realization of perovskite-based phononic transducers, including those for chiral phonons.